OpenStudy (anonymous):

2,401^2n=343^n+2 someone hlep me slove this asap lol

OpenStudy (helder_edwin):

\[ \large 2401^{2n}=343^{n+2} \] right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

\[ \large 343=7^3 \] \[ \large 2401=7^4 \]

OpenStudy (helder_edwin):

so \[ \large 2401^{2n}=343^{n+2} \] \[ \large (7^4)^{2n}=(7^3)^{n+2} \] \[ \large 7^{8n}=7^{3n+6} \]

OpenStudy (helder_edwin):

do u get this so far?

OpenStudy (anonymous):

yes lol

OpenStudy (helder_edwin):

can u finish?

OpenStudy (anonymous):

No i suck at math :/

OpenStudy (anonymous):

i wish i didnt

OpenStudy (helder_edwin):

the next step is \[ \large 8n=3n+6 \] \[ \large 8n-3n=6 \] \[ \large 5n=6 \] \[ \large n=6/5 \]

OpenStudy (anonymous):

thank you sooooooo much!

OpenStudy (helder_edwin):

u r w but, did u understand?

OpenStudy (anonymous):

yews i did. I just never knew how to do a problem like thaqt and noe i do i wrote it all down too! :P haha